# Light Dirac neutrino portal dark matter with gauged $U(1)_{B-L}$ symmetry

### 11 Dec 2023 v1

@@ I Introduction
In this work, we consider a scenario where the origin of DM is related to the Dirac nature of light neutrinos, known as the light Dirac neutrino portal DM Biswas:2021kio ; Biswas:2022vkq scenario. In such a setup, light Dirac neutrinos take the role of mediating the interactions between DM and the SM bath. In Biswas:2021kio and Biswas:2022vkq , the DM was assumed to be of WIMP and FIMP type respectively. In addition to linking the origin of neutrino mass and nature with DM, this also offers additional discovery prospects due to right chiral part of Dirac neutrinos, contributing to the effective relativistic degrees of freedom ${\rm N_{eff}}$. Measurement related to the cosmic microwave background (CMB) puts tight constraints ${\rm N_{eff}=2.99^{+0.34}_{-0.33}}$ at $2\sigma$ or $95\%$ CL including baryon acoustic oscillation (BAO) data Planck:2018vyg . Similar bound also exists from big bang nucleosynthesis (BBN) $3.42.3<{\rm N}_{\rm eff}<$ 3.4 at $95\%$ CL Cyburt:2015mya . Both of these cosmological bounds are consistent with the SM predictions ${\rm N^{SM}_{eff}}=3.045$ Mangano:2005cc ; Grohs:2015tfy ; deSalas:2016ztq  [^3] A very recent paper Cielo:2023bqp reports $N_{\rm eff}^{\rm SM}=3.043$ by incorporating next-to-leading order correction to $e^{+}e^{-}\leftrightarrow\nu_{L}\bar{\nu}_{L}$ interactions along with finite temperature QED corrections to the electromagnetic plasma density and effect of neutrino oscillations. . Future CMB experiment CMB Stage IV (CMB-S4) is expected reach a much better sensitivity of $\Delta{\rm N}_{\rm eff}={\rm N}_{\rm eff}-{\rm N}^{\rm SM}_{\rm eff}=0.06$ Abazajian:2019eic , taking it closer to the SM prediction. Thus, Dirac neutrino scenarios can be probed in future CMB experiments if right handed neutrinos (RHN) can be sufficiently produced in the early universe via thermal or non-thermal processes. Light Dirac neutrino models often lead to enhanced $\Delta{\rm N}_{\rm eff}$, some recent works on which can be found in Abazajian:2019oqj ; FileviezPerez:2019cyn ; Nanda:2019nqy ; Han:2020oet ; Luo:2020sho ; Borah:2020boy ; Adshead:2020ekg ; Luo:2020fdt ; Mahanta:2021plx ; Du:2021idh ; Biswas:2021kio ; Borah:2022obi ; Borah:2022qln ; Li:2022yna ; Biswas:2022fga ; Borah:2023dhk ; Borah:2022enh ; Das:2023oph . In earlier works on light Dirac neutrino portal DM, discrete symmetry like $Z_{4}$ was considered to get the desired couplings, mass terms as well as the stability of DM. Here, we consider a UV completion with a gauged $B-L$ framework. The right handed neutrinos are naturally part of the model providing the minimal anomaly-free setup. While RHNs are thermally produced due to gauged $B-L$ interactions, DM can be of WIMP or FIMP type depending upon Dirac neutrino portal couplings. The model not only gives rise to the desired DM phenomenology with observable $\Delta{\rm N}_{\rm eff}$, but also leads to new constraints in the gauged $B-L$ parameter space not obtained previously.
@@ II The Model
Gauged $B-L$ extension of the SM Davidson:1978pm ; Mohapatra:1980qe ; Marshak:1979fm ; Masiero:1982fi ; Mohapatra:1982xz ; Buchmuller:1991ce has been a popular BSM framework studied in the context of neutrino mass among others. The three right handed neutrinos $\nu_{R}$ having $B-L$ charge -1 each not only keep the model anomaly free but also lead to massive Dirac neutrinos in combination with $\nu_{L}$ after electroweak symmetry breaking. A singlet fermion $\psi$ is considered to be the DM candidate while two singlet scalars $\phi_{1}$ and $\phi_{2}$ with non-zero $B-L$ charges help in realising light Dirac neutrino portal DM and spontaneous $B-L$ symmetry breaking respectively. The relevant particle content is shown in table 1.
@@
Considering the the free parameters in the model to be $\mu_{1},\mu_{2},\lambda_{1},\lambda_{2},\lambda_{H\phi_{1}},g_{BL},y_{\phi_{1}},m_{\psi}$, the physical masses of $\phi_{1},\phi_{2}$ and $Z^{\prime}$ (the $U(1)_{B-L}$ gauge boson) can be written as
@@
where $v\left(=\sqrt{\frac{\mu^{2}}{\lambda}}\right)$ and $v_{2}\left(=\sqrt{\frac{\mu_{2}^{2}}{\lambda_{2}}}\right)$ are the VEVs of the neutral component of the SM Higgs $\Phi$ and $\phi_{2}$ respectively. As a result, the parameters $\mu_{1},\mu_{2}$ and $\lambda_{2}$ can be traded for the parameters $m_{\phi_{1}},m_{\phi_{2}}$ and $m_{Z^{\prime}}$. This leads to the free parameters as $m_{\phi_{1}},m_{\phi_{2}},m_{Z^{\prime}},\lambda_{1},\lambda_{H\phi_{1}},g_{BL},y_{\phi_{1}},m_{\psi}$, out of which the relevant parameters for the phenomenology to be discussed are $m_{\phi_{1}},m_{Z^{\prime}},m_{\psi},\lambda_{H\phi_{1}},g_{BL},y_{\phi_{1}}$.
@@ III FIMP type dark matter
In this scenario, the Yukawa coupling among $\nu_{R}$, DM and singlet scalar is small $y_{\phi_{1}}<10^{-7}$ keeping the associated processes out-of-equilibrium throughout. While DM never thermalises, the RHNs can thermalise with the SM bath by virtue of $U(1)_{B-L}$ gauge interactions. Similar to the active neutrinos, the RHNs also decouple from the bath at a temperature $T_{\nu_{R}}$ and contributes to the effective relativistic degrees of freedom defined as
$N_{\rm eff}\equiv\frac{8}{7}\left(\frac{11}{4}\right)^{4/3}\left(\frac{\rho_{\rm rad}-\rho_{\gamma}}{\rho_{\gamma}}\right),$
where $\rho_{\rm rad}$ is the net radiation content of the universe. As mentioned earlier, the SM prediction is $N^{\rm SM}_{\rm eff}=3.045$ Mangano:2005cc ; Grohs:2015tfy ; deSalas:2016ztq . In our model, $\Delta N_{\rm eff}$ can be estimated by finding the decoupling temperature $T_{\nu_{R}}$ of right handed neutrinos ($\nu_{R}$) using
@@
Figure 1: $\rm\Delta N_{\rm eff}$ as a function of $m_{Z^{\prime}}$ and $g_{BL}$.
@@
After studying the evolution of comoving densities for different benchmark points, we perform a numerical scan over key model parameters to find out the parameter space consistent with $\Delta N_{\rm eff}$ and DM properties. We have kept $m_{\phi}$ fixed at $1000$ GeV and $y_{\phi_{1}}$ at $10^{-10}$. The rest of the parameters are varied in the following range:
@@
While we have incorporated the constraints from cosmological observations on $\Delta N_{\rm eff}$ and DM relic abundance, there can be strong constraints on light dark matter from astrophysical structure formation. Such bounds can be imposed on a particular DM scenario by calculating the free-streaming length (FSL) of DM. While hot DM is already ruled out, warm DM with FSL $\lambda_{\rm FSL}<0.1$ Mpc is still allowed, and can be favourable over cold DM of FSL $\lambda_{\rm FSL}<0.01$ Mpc due to the small-scale structure problems associated with the latter Drewes:2016upu . Dark matter free-streaming length can be estimated from matter power spectrum inferred from the Lyman-$\alpha$ forest data Croft:2000hs ; Kim:2003qt ; Viel:2005qj ; Irsic:2017ixq and also from Quasar data Hsueh:2019ynk . Such estimates are also supported by theoretical and simulation based results Colombi:1995ze ; Boyarsky:2008xj ; deVega:2009ku ; Schneider:2011yu . For some recent discussions on structure formation constraints on DM production mechanisms, please see Merle:2013wta ; Decant:2021mhj ; Ballesteros:2020adh and references therein. The detailed calculations related to FSL of DM are given in appendix A.
As shown in appendix A, the FSL of DM depends primarily on the production temperature, DM mass and the production mechanism of DM or distribution function of DM. A higher production temperature gives a smaller FSL due to high momentum redshift making DM non-relativistic earlier and vice-versa. A different production mechanism also gives different FSL. In our case, the DM is produced due to decay of a frozen out scalar. Fig. 4 shows the average velocity of DM for four different benchmark parameters. The parameters are shown in table 2 along with total $\Delta N_{\rm eff}$ and FSL. The benchmark point (BP) II corresponds to a smaller $\phi_{1}$ mass than BP I. Hence, the freeze-out abundance of $\phi_{1}$ for BP II is smaller than that of BP I. This implies that a larger $\psi$ mass is required to satisfy DM abundance. As a result, we obtain a smaller FSL for BP II. However, for both BP I and BP II, the computed FSL keeps DM in hot DM category and hence ruled out from structure formation constraints. The FSL can be reduced by increasing the production temperature as well as increasing DM mass. BP III and BP IV have a larger $y_{\phi_{1}}$ giving a higher production temperature. Similarly it has a larger $\lambda_{H\phi_{1}}$ coupling, giving smaller $\phi_{1}$ freeze-out abundance. Hence for both BP III and BP IV, DM masses can be large while being consistent with relic abundance criteria. Combining these two effects, we get warm dark matter for BP III and BP IV. From the resulting on $\Delta N_{\rm eff}$, it can be seen that both BP III and BP IV remain within the sensitivity of future CMB experiments like CMB-S4.
@@
Fig. 5 shows the parameter space in $m_{Z^{\prime}}$ versus $g_{BL}$ parameter space for the FIMP scenario. The left, middle and right panel plots correspond to dark matter mass of $400$ keV, $500$ keV and $1300$ keV respectively. The other relevant parameters are fixed as $m_{\phi_{1}}=500$ GeV, $\lambda_{H\phi_{1}}=10^{-5}$, $y_{\phi_{1}}=10^{-9}$. For the region below the solid brown line in each of these plots, $U(1)_{B-L}$-portal coupling does not play any role in FIMP DM production. In that region, the freeze-out abundance of $\phi_{1}$, responsible for DM production, is determined by the Higgs-portal coupling $\lambda_{H\phi_{1}}$. In other words, the cross section $\langle{\sigma v}\rangle_{{}_{\phi_{1}\phi_{1}^{\dagger{}}\to SM\overline{SM}}}$ dominates over $\langle{\sigma v}\rangle_{{}_{\phi_{1}\phi_{1}^{\dagger{}}\to Z^{\prime}Z^{\prime}}}$. In the region above the solid brown line, the freeze-out abundance of $\phi_{1}$ is determined by the $U(1)_{B-L}$-portal coupling. The dashed line in the left and middle panel plots, separating the pink and cyan shaded regions, represent the parameter space satisfying correct DM relic for $m_{\psi}=400$ keV and $m_{\psi}=500$ keV respectively. However, for DM mass $m_{\psi}=400$ keV, the free-streaming length turns out to be very large $\sim 0.1$ Mpc, keeping it in hot DM ballpark and hence ruled out. In the middle panel plot, due to a higher DM mass $m_{\psi}=500$ keV, the dashed line corresponds to an intermediate FSL keeping it in the warm DM regime which is still allowed by structure formation constraints. In the region below the dashed line, the freeze-out abundance of $\phi_{1}$ is more than that for the dashed line, leading to a larger FIMP DM abundance $\Omega_{\rm DM}h^{2}>0.12$, produced from $\phi_{1}$ decay. On the other hand, the region above the dashed line have $\Omega_{\rm DM}h^{2}<0.12$ due to smaller freeze-out abundance of $\phi_{1}$. This is understood from the fact that a larger $U(1)_{B-L}$-portal interaction leads to smaller freeze-out abundance of $\phi_{1}$ and vice versa. While the total $\Delta N_{\rm eff}$ can have both thermal and non-thermal contributions, for the choice of parameters in Fig. 5, the non-thermal contribution is suppressed compared to the thermal contribution. The magenta solid line corresponds to $\Delta N_{\rm eff}=0.28$. The region above this line has $\Delta N_{\rm eff}>0.28$ and hence ruled out from Planck 2018 limits. In the right panel plot, the dashed line corresponds to relic satisfying region for DM mass $m_{\psi}=1300$ keV. The region below the dashed line correspond to over-abundance like before. However, the region above the dashed line is no longer under-abundant, but satisfies the current DM abundance. For this region of parameter space and chosen DM mass, the production of FIMP DM while $\phi_{1}$ is in equilibrium dominates compared to the production after $\phi_{1}$ freeze-out. This can also be understood from the evolution plots shown in Fig. 2. From the left panel plot of Fig. 2, we see that there is a non-zero yield of DM $Y_{\psi}$ from the decay of $\phi_{1}$ when the latter is in equilibrium. Additionally, the freeze-out abundance of $\phi_{1}$ ($Y_{\phi_{1}}(x\rightarrow\infty)$) decreases as we increase $g_{BL}$. Combining these two, we can have a situation where for sufficiently large $U(1)_{B-L}$-portal couplings, $Y_{\psi}$ from the decay of $\phi_{1}$ in equilibrium is larger compared to the post freeze-out production. Hence, $U(1)_{B-L}$-portal couplings do not determine the comoving abundance of $Y_{\psi}$ and as a result, the region above the dashed line satisfies the correct DM relic in the right panel plot of Fig. 5. In all the plots, the region below the magenta solid line has $0.140.28>\Delta N_{\rm eff}>$ 0.14 keeping it within reach of future CMB experiments like CMB-S4. The blue shaded region towards the upper right corner indicates the region ruled out from the large hadron collider (LHC) bounds, for specifically from ATLAS experiments at 13 TeV centre of mass energy Aaboud:2017buh ; Aad:2019fac ; ATLAS:2017fih . Similar bounds also exist from the CMS experiment at the LHC Sirunyan:2018exx . A relatively weaker bound exists from the large electron positron (LEP) collider disfavouring the region $m_{Z^{\prime}}/g_{BL}<7$ TeV Carena:2004xs ; Cacciapaglia:2006pk . The parameter space shown in Fig. 5 already satisfies the LEP bound.
@@
$Y_{\psi}(x\rightarrow\infty)$ $=$ $\frac{135\,}{16\pi\times 1.66\times 8\pi^{3}g^{s}_{*}\sqrt{g^{\rho}_{*}}}M_{\rm pl}\,\frac{y^{2}_{\phi_{1}}}{m_{\phi}}\,,$
$\widetilde{Y}(x\rightarrow\infty)$ $=$ $\frac{675\,}{32\pi\times 1.66\times 8\pi^{3}g^{s}_{*}\sqrt{g^{\rho}_{*}}}\left(\frac{45}{2\pi^{2}\,g^{s}_{*}}\right)^{1/3}M_{\rm pl}\,\frac{y^{2}_{\phi_{1}}}{m_{\phi}}\,,$ (15)
@@
where $s^{0}$ and $\rho^{0}_{c}$ denote the entropy density and critical energy density, respectively, at the present epoch. $\rho_{\nu_{L}}$ denotes energy density of one species of SM neutrino and $h$ represents $H_{0}/100$ with $H_{0}$ being the current expansion rate of the universe. Here, $m_{\phi_{1}},y_{\phi_{1}}$ and $m_{\psi}$ determine the DM abundance and $\Delta N^{\rm non-th}_{\rm eff}$ is determined by $m_{\phi_{1}}$ and $y_{\phi_{1}}$ only. The thermal contribution to $N_{\rm eff}$ is determined by the parameters $m_{Z^{\prime}}$ and $g_{BL}$. When DM abundance is satisfied, it turns out that the non-thermal contribution to $\Delta N_{\rm eff}$ is way below the minimum thermal contribution. So, the total $\Delta N_{\rm eff}$ is only determined by the thermal contribution (shown in Fig. 1).

### 09 Apr 2024 v2

@@ I Introduction
In this work, we consider a scenario where the origin of DM is related to the Dirac nature of light neutrinos, known as the light Dirac neutrino portal DM Biswas:2021kio ; Biswas:2022vkq scenario [^3] Neutrino portal DM has been studied in several earlier works where either SM light neutrino or heavy neutrinos were considered as the portal Falkowski:2009yz ; Macias:2015cna ; Batell:2017rol ; Batell:2017cmf ; Bandyopadhyay:2018qcv ; Chianese:2018dsz ; Blennow:2019fhy ; Lamprea:2019qet ; Chianese:2019epo ; Bandyopadhyay:2020qpn ; Hall:2019rld ; Berlin:2018ztp .. In such a setup, light Dirac neutrinos take the role of mediating the interactions between DM and the SM bath. In Biswas:2021kio and Biswas:2022vkq , the DM was assumed to be of WIMP and FIMP type respectively. In addition to linking the origin of neutrino mass and nature with DM, this also offers additional discovery prospects due to right chiral part of Dirac neutrinos, contributing to the effective relativistic degrees of freedom ${\rm N_{eff}}$. Measurement related to the cosmic microwave background (CMB) puts tight constraints ${\rm N_{eff}=2.99^{+0.34}_{-0.33}}$ at $2\sigma$ or $95\%$ CL including baryon acoustic oscillation (BAO) data Planck:2018vyg . Similar bound also exists from big bang nucleosynthesis (BBN) $3.42.3<{\rm N}_{\rm eff}<$ 3.4 at $95\%$ CL Cyburt:2015mya . Both of these cosmological bounds are consistent with the SM predictions ${\rm N^{SM}_{eff}}=3.045$ Mangano:2005cc ; Grohs:2015tfy ; deSalas:2016ztq  [^4] A very recent paper Cielo:2023bqp reports $N_{\rm eff}^{\rm SM}=3.043$ by incorporating next-to-leading order correction to $e^{+}e^{-}\leftrightarrow\nu_{L}\bar{\nu}_{L}$ interactions along with finite temperature QED corrections to the electromagnetic plasma density and effect of neutrino oscillations. . Future CMB experiment CMB Stage IV (CMB-S4) is expected reach a much better sensitivity of $\Delta{\rm N}_{\rm eff}={\rm N}_{\rm eff}-{\rm N}^{\rm SM}_{\rm eff}=0.06$ Abazajian:2019eic , taking it closer to the SM prediction. Thus, Dirac neutrino scenarios can be probed in future CMB experiments if right handed neutrinos (RHN) can be sufficiently produced in the early universe via thermal or non-thermal processes. Light Dirac neutrino models often lead to enhanced $\Delta{\rm N}_{\rm eff}$, some recent works on which can be found in Abazajian:2019oqj ; FileviezPerez:2019cyn ; Nanda:2019nqy ; Han:2020oet ; Luo:2020sho ; Borah:2020boy ; Adshead:2020ekg ; Luo:2020fdt ; Mahanta:2021plx ; Du:2021idh ; Biswas:2021kio ; Borah:2022obi ; Borah:2022qln ; Li:2022yna ; Biswas:2022fga ; Adshead:2022ovo ; Borah:2023dhk ; Borah:2022enh ; Das:2023oph . In earlier works on light Dirac neutrino portal DM, discrete symmetry like $Z_{4}$ was considered to get the desired couplings, mass terms as well as the stability of DM. Here, we consider a UV completion with a gauged $B-L$ framework. The right handed neutrinos are naturally part of the model providing the minimal anomaly-free setup. While RHNs are thermally produced due to gauged $B-L$ interactions, DM can be of WIMP or FIMP type depending upon Dirac neutrino portal couplings. The model not only gives rise to the desired DM phenomenology with observable $\Delta{\rm N}_{\rm eff}$, but also leads to new constraints in the gauged $B-L$ parameter space not obtained previously.
@@ II The Model
Gauged $B-L$ extension of the SM Davidson:1978pm ; Mohapatra:1980qe ; Marshak:1979fm ; Masiero:1982fi ; Mohapatra:1982xz ; Buchmuller:1991ce has been a popular BSM framework studied in the context of neutrino mass among others. The three right handed neutrinos $\nu_{R}$ having $B-L$ charge -1 each not only keep the model anomaly free [^5] See appendix A for other solutions to anomaly cancellation conditions. but also lead to massive Dirac neutrinos in combination with $\nu_{L}$ after electroweak symmetry breaking. A singlet fermion $\psi$ is considered to be the DM candidate while two singlet scalars $\phi_{1}$ and $\phi_{2}$ with non-zero $B-L$ charges help in realising light Dirac neutrino portal DM and spontaneous $B-L$ symmetry breaking respectively. The relevant particle content is shown in table 1.
@@
Considering the the free parameters in the model to be $\mu_{1},\mu_{2},\lambda_{1},\lambda_{2},\lambda_{H\phi_{1}},{\lambda^{\prime}_{\phi_{1}\phi_{2}}},g_{BL},y_{\phi_{1}},m_{\psi}$, the physical masses of $\phi_{1},\phi_{2}$ and $Z^{\prime}$ (the $U(1)_{B-L}$ gauge boson) can be written as
@@
where $v\left(=\sqrt{\frac{\mu^{2}}{\lambda}}\right)$ and $v_{2}\left(=\sqrt{\frac{\mu_{2}^{2}}{\lambda_{2}}}\right)$ are the VEVs of the neutral component of the SM Higgs $\Phi$ and $\phi_{2}$ respectively. As a result, the parameters $\mu_{1},\mu_{2}$ and $\lambda_{2}$ can be traded for the parameters $m_{\phi_{1}},m_{\phi_{2}}$ and $m_{Z^{\prime}}$. This leads to the free parameters as $m_{\phi_{1}},m_{\phi_{2}},m_{Z^{\prime}},\lambda_{1},\lambda_{H\phi_{1}},{\lambda^{\prime}_{\phi_{1}\phi_{2}}},g_{BL},y_{\phi_{1}},m_{\psi}$, out of which the relevant parameters for the phenomenology to be discussed are 666 The scalar $\phi_{2}$ is not expected to play any role in dark matter analysis. First of all, DM does not couple directly to $\phi_{2}$. Also, $\phi_{2}$ is heavier than the rest of the particles and hence does not appear in final states. $m_{\phi_{1}},m_{Z^{\prime}},m_{\psi},\lambda_{H\phi_{1}},g_{BL},y_{\phi_{1}}$.
@@ III FIMP type dark matter
In this scenario, the Yukawa coupling among $\nu_{R}$, DM and singlet scalar is small $y_{\phi_{1}}<10^{-7}$ keeping the associated processes out-of-equilibrium throughout. While DM never thermalises, the RHNs can thermalise with the SM bath by virtue of $U(1)_{B-L}$ gauge interactions. Similar to the active neutrinos, the RHNs also contributes to the effective relativistic degrees of freedom defined as
$N_{\rm eff}\equiv\frac{8}{7}\left(\frac{11}{4}\right)^{4/3}\left(\frac{\rho_{\rm rad}-\rho_{\gamma}}{\rho_{\gamma}}\right),$ (8)
where $\rho_{\rm rad}$ is the net radiation content of the universe  [^7] Due to the presence of three relativistic RHNs, change in $N_{\rm eff}$ can be written as $\Delta N_{\rm eff}=3\left.{\frac{\rho_{\nu_{R}}}{\rho_{\nu_{L}}}}\right|_{\rm CMB}=3\left.{\frac{\rho_{\nu_{R}}}{\rho_{\nu_{L}}}}\right|_{\rm 10MeV}=3\left.{\left(\frac{T_{\nu_{R}}}{T_{\nu_{L}}}\right)^{4}}\right|_{\rm 10MeV}$. Here $\rho_{\nu_{L}}$ denotes energy density of one species of SM neutrinos. For non-thermal RHNs, $\rho_{\nu_{R}}$ should be calculated by solving appropriate Boltzmann equation whereas for thermal RHNs, $\rho_{\nu_{R}}$ can be written in terms of its decoupling temperature.. As mentioned earlier, the SM prediction is $N^{\rm SM}_{\rm eff}=3.045$ Mangano:2005cc ; Grohs:2015tfy ; deSalas:2016ztq . In our model for thermal right handed neutrinos ($\nu_{R}$), $\Delta N_{\rm eff}$ can be estimated by finding the decoupling temperature $T_{\nu_{R}}$ of $\nu_{R}$ using
@@
Figure 1: $\rm\Delta N^{\rm th}_{\rm eff}$ as a function of $m_{Z^{\prime}}$ and $g_{BL}$.
@@
After studying the evolution of comoving densities for different benchmark points, we perform a numerical scan over key model parameters to find out the parameter space consistent with $\Delta N_{\rm eff}$ and DM properties. We have kept $m_{\phi_{1}}$ fixed at $1000$ GeV and $y_{\phi_{1}}$ at $10^{-10}$. The rest of the parameters are varied in the following range:
@@
While we have incorporated the constraints from cosmological observations on $\Delta N_{\rm eff}$ and DM relic abundance, there can be strong constraints on light dark matter from astrophysical structure formation. Such bounds can be imposed on a particular DM scenario by calculating the free-streaming length (FSL) of DM. While hot DM is already ruled out, warm DM with FSL $\lambda_{\rm FSL}<0.1$ Mpc is still allowed, and can be favourable over cold DM of FSL $\lambda_{\rm FSL}<0.01$ Mpc due to the small-scale structure problems associated with the latter Drewes:2016upu . Dark matter free-streaming length can be estimated from matter power spectrum inferred from the Lyman-$\alpha$ forest data Croft:2000hs ; Kim:2003qt ; Viel:2005qj ; Irsic:2017ixq and also from Quasar data Hsueh:2019ynk . Such estimates are also supported by theoretical and simulation based results Colombi:1995ze ; Boyarsky:2008xj ; deVega:2009ku ; Schneider:2011yu . For some recent discussions on structure formation constraints on DM production mechanisms, please see Merle:2013wta ; Decant:2021mhj ; Ballesteros:2020adh and references therein. The detailed calculations related to FSL of DM are given in appendix B.
As shown in appendix B, the FSL of DM depends primarily on the production temperature, DM mass and the production mechanism of DM or distribution function of DM. A higher production temperature gives a smaller FSL due to high momentum redshift making DM non-relativistic earlier and vice-versa. A different production mechanism also gives different FSL. In our case, the DM is produced due to decay of a frozen out scalar. Fig. 4 shows the average velocity of DM for four different benchmark parameters. The parameters are shown in table 2 along with total $\Delta N_{\rm eff}$ and FSL. The benchmark point (BP) II corresponds to a smaller $\phi_{1}$ mass than BP I. Hence, the freeze-out abundance of $\phi_{1}$ for BP II is smaller than that of BP I. This implies that a larger $\psi$ mass is required to satisfy DM abundance. As a result, we obtain a smaller FSL for BP II. However, for both BP I and BP II, the computed FSL keeps DM in hot DM category and hence ruled out from structure formation constraints. The FSL can be reduced by increasing the production temperature as well as increasing DM mass. BP III and BP IV have a larger $y_{\phi_{1}}$ giving a higher production temperature. Similarly it has a larger $\lambda_{H\phi_{1}}$ coupling, giving smaller $\phi_{1}$ freeze-out abundance. Hence for both BP III and BP IV, DM masses can be large while being consistent with relic abundance criteria. Combining these two effects, we get warm dark matter for BP III and BP IV. From the resulting on $\Delta N_{\rm eff}$, it can be seen that both BP III and BP IV remain within the sensitivity of future CMB experiments like CMB-S4.
@@
Fig. 5 shows the parameter space in $m_{Z^{\prime}}$ versus $g_{BL}$ parameter space for the FIMP scenario. The left, middle and right panel plots correspond to dark matter mass of $400$ keV, $500$ keV and $1300$ keV respectively. The other relevant parameters are fixed as $m_{\phi_{1}}=500$ GeV, $\lambda_{H\phi_{1}}=10^{-5}$, $y_{\phi_{1}}=10^{-9}$. For the region below the solid brown line in each of these plots, $U(1)_{B-L}$-portal coupling does not play any role in FIMP DM production. In that region, the freeze-out abundance of $\phi_{1}$, responsible for DM production, is determined by the Higgs-portal coupling $\lambda_{H\phi_{1}}$. In other words, the cross section $\langle{\sigma v}\rangle_{{}_{\phi_{1}\phi_{1}^{\dagger{}}\to SM\overline{SM}}}$ dominates over $\langle{\sigma v}\rangle_{{}_{\phi_{1}\phi_{1}^{\dagger{}}\to Z^{\prime}Z^{\prime}}}$. In the region above the solid brown line, the freeze-out abundance of $\phi_{1}$ is determined by the $U(1)_{B-L}$-portal coupling. The dashed line in the left and middle panel plots, separating the pink and cyan shaded regions, represent the parameter space satisfying correct DM relic for $m_{\psi}=400$ keV and $m_{\psi}=500$ keV respectively. However, for DM mass $m_{\psi}=400$ keV, the free-streaming length turns out to be very large $\sim 0.1$ Mpc, keeping it in hot DM ballpark and hence ruled out. In the middle panel plot, due to a higher DM mass $m_{\psi}=500$ keV, the dashed line corresponds to an intermediate FSL keeping it in the warm DM regime which is still allowed by structure formation constraints. In the region below the dashed line, the freeze-out abundance of $\phi_{1}$ is more than that for the dashed line, leading to a larger FIMP DM abundance $\Omega_{\rm DM}h^{2}>0.12$, produced from $\phi_{1}$ decay. On the other hand, the region above the dashed line have $\Omega_{\rm DM}h^{2}<0.12$ due to smaller freeze-out abundance of $\phi_{1}$. This is understood from the fact that a larger $U(1)_{B-L}$-portal interaction leads to smaller freeze-out abundance of $\phi_{1}$ and vice versa. While the total $\Delta N_{\rm eff}$ can have both thermal and non-thermal contributions, for the choice of parameters in Fig. 5, the non-thermal contribution is suppressed compared to the thermal contribution. The magenta solid line corresponds to $\Delta N_{\rm eff}=0.28$. The region above this line has $\Delta N_{\rm eff}>0.28$ and hence ruled out from Planck 2018 limits. In the right panel plot, the dashed line corresponds to relic satisfying region for DM mass $m_{\psi}=1300$ keV. The region below the dashed line correspond to over-abundance like before. However, the region above the dashed line is no longer under-abundant, but satisfies the current DM abundance. For this region of parameter space and chosen DM mass, the production of FIMP DM while $\phi_{1}$ is in equilibrium dominates compared to the production after $\phi_{1}$ freeze-out. This can also be understood from the evolution plots shown in Fig. 2. From the left panel plot of Fig. 2, we see that there is a non-zero yield of DM $Y_{\psi}$ from the decay of $\phi_{1}$ when the latter is in equilibrium. Additionally, the freeze-out abundance of $\phi_{1}$ ($Y_{\phi_{1}}(x\rightarrow\infty)$) decreases as we increase $g_{BL}$. Combining these two, we can have a situation where for sufficiently large $U(1)_{B-L}$-portal couplings, $Y_{\psi}$ from the decay of $\phi_{1}$ in equilibrium is larger compared to the post freeze-out production. Hence, $U(1)_{B-L}$-portal couplings do not determine the comoving abundance of $Y_{\psi}$ and as a result, the region above the dashed line satisfies the correct DM relic in the right panel plot of Fig. 5. In all the plots, the region below the magenta solid line has $0.140.28>\Delta N_{\rm eff}>$ 0.14 keeping it within reach of future CMB experiments like CMB-S4. The blue shaded region towards the upper right corner indicates the region ruled out from the large hadron collider (LHC) bounds, for specifically from ATLAS experiments at 13 TeV centre of mass energy ATLAS:2019erb ; ATLAS:2017fih . Similar bounds also exist from the CMS experiment at the LHC Sirunyan:2018exx . The details of the LHC bound on $g_{BL}-m_{Z^{\prime}}$ plane is given in appendix C. A relatively weaker bound exists from the large electron positron (LEP) collider disfavouring the region $m_{Z^{\prime}}/g_{BL}<7$ TeV Carena:2004xs ; Cacciapaglia:2006pk . The parameter space shown in Fig. 5 already satisfies the LEP bound.
@@
$Y_{\psi}(x\rightarrow\infty)$ $=$ $\frac{135\,}{16\pi\times 1.66\times 8\pi^{3}g^{s}_{*}\sqrt{g^{\rho}_{*}}}M_{\rm pl}\,\frac{y^{2}_{\phi_{1}}}{m_{\phi_{1}}}\,,$
$\widetilde{Y}(x\rightarrow\infty)$ $=$ $\frac{675\,}{32\pi\times 1.66\times 8\pi^{3}g^{s}_{*}\sqrt{g^{\rho}_{*}}}\left(\frac{45}{2\pi^{2}\,g^{s}_{*}}\right)^{1/3}M_{\rm pl}\,\frac{y^{2}_{\phi_{1}}}{m_{\phi_{1}}}\,,$ (16)
@@
where $s^{0}$ and $\rho^{0}_{c}$ denote the entropy density and critical energy density, respectively, at the present epoch. $\rho_{\nu_{L}}$ denotes energy density of one species of SM neutrino and $h$ represents $H_{0}/100$ with $H_{0}$ being the current expansion rate of the universe. Here, $m_{\phi_{1}},y_{\phi_{1}}$ and $m_{\psi}$ determine the DM abundance and $\Delta N^{\rm non-th}_{\rm eff}$ is determined by $m_{\phi_{1}}$ and $y_{\phi_{1}}$ only. The thermal contribution to $\Delta N_{\rm eff}$ is determined by the parameters $m_{Z^{\prime}}$ and $g_{BL}$. When DM abundance is satisfied, it turns out that the non-thermal contribution to $\Delta N_{\rm eff}$ is way below the minimum thermal contribution. So, the total $\Delta N_{\rm eff}$ is only determined by the thermal contribution (shown in Fig. 1).